OpenStudy (el_arrow):

Consider a pitcher throwing a ball. If he is on a tower so that the ball is 5m off the ground when thrown horizontally, and the ball lands 18m downrange, what is the speed of the ball?

OpenStudy (el_arrow):

someone please help me

OpenStudy (el_arrow):

@ganeshie8

OpenStudy (el_arrow):

@freckles

OpenStudy (anonymous):

You need two steps to get to the solution. Step #1: calculate the time taken for the ball to hit the ground This is only dependent on the vertical motion of the ball. So you can find the time by assuming the ball is just falling under the influence of gravity:\[h = \frac{1}{2}gt^{2}\] Step #2: calculate the horizontal speed of the ball By understanding that with no air resistance the ball will move at the same velocity for its flight, you can write that its distance traveled is simply the product of velocity and time.\[x = vt\]

OpenStudy (anonymous):

El

OpenStudy (el_arrow):

is the equation to find t this one t= sqrt(.5*g)/h

OpenStudy (el_arrow):

@Abhisar can you help me real quick

OpenStudy (abhisar):



OpenStudy (el_arrow):

okay

OpenStudy (el_arrow):

so do i use Pythagorean theoream?

OpenStudy (abhisar):

Wait.. (-:

OpenStudy (abhisar):

ok, You need to find out the horizontal speed by which the ball was thrown..right?

OpenStudy (el_arrow):

yes

OpenStudy (abhisar):

So to find that we would have to first find out the time of flight of the ball.

OpenStudy (el_arrow):

yes

OpenStudy (abhisar):

The ball when thrown had two motions, one in the horizontal direction and second in the vertical direction. We can find out the time by using its vertical motion. We can use the equation \(\tt S=ut+\frac{1}{2}a\times t^2\) to find out the time.

OpenStudy (abhisar):

Calculate the time, what would it be?

OpenStudy (el_arrow):

is that a kinematic equation?

OpenStudy (abhisar):

Yes (-:

OpenStudy (el_arrow):

so we need the t on the left side right/

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (el_arrow):

would it be t=1/2a / s-u?

OpenStudy (abhisar):

Just plug in the values, s will be 5, u will be 0 since ball was thrown horizontally and a will be 10 (acceleration due to gravity).

OpenStudy (abhisar):

So, the equation will be like \(\tt 5 = 0\times t + \frac{1}{2}\times 10\times t^2\), find t

OpenStudy (el_arrow):

t = sqrt(1/5)

OpenStudy (abhisar):

Nopes.....

OpenStudy (el_arrow):

oh crap forgot the 1/2 its sqrt of 1

OpenStudy (el_arrow):

right?

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (abhisar):

Now the second and final part comes in. You can use the same equation to calculate the horizontal speed. This time we will consider ball's horizontal motion. Take s=18, t= 1 and a=0 (there is no horizontal acceleration). Plugging in the values we get, \(\tt 18 = u\times 1+\frac{1}{2}\times 0 \times 1^2\), find the value of u

OpenStudy (el_arrow):

u= 18

OpenStudy (el_arrow):

right?

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (abhisar):

That's the answer (-:

OpenStudy (el_arrow):

thank you so much

OpenStudy (abhisar):

Welcome (-: